OpenStudy (anonymous):

Evaluate 7c - 2bc, when b = 2 and c = 3. A. 114 B. 9 C. 2 D. 33

OpenStudy (anonymous):

Your answer is B. 7(3)-2(2)(3)=21-12=9!

OpenStudy (anonymous):

7*3-2*2*3=114 a.114

OpenStudy (anonymous):

B: 9